OpenStudy (anonymous):

How fast would an unladen sparrow (28.3 g) have to travel in order to have the same momentum as a baseball (0.145 kg) that is thrown at 40.0 m/s (~ 89.5 mph).

OpenStudy (michele_laino):

here, we have to apply this formula: \[\large {v_s} = \frac{{{m_b}{v_b}}}{{{m_s}}} = \frac{{0.145 \cdot 40}}{{0.0283}} = ...?\] where \(\large m_s, v_s\) are the mass and the speed of the sparrow, whereas \( \large m_b, v_b\) are the corresponding quantities of the ball